OpenStudy (anonymous):

In a thunderstorm at 20.0 C Amanda is able to see a lightning strike and hears the thunderclap about 3.0 s afterwards. How far was the lightning strike from Amanda?

OpenStudy (anonymous):

@roadjester

OpenStudy (roadjester):

Calculate the speed of sound at a specific temperatures. There's an equation for it that I can't remember but it's in a square root.

OpenStudy (anonymous):

I found this online (331 + (0.6 *20) = 343m/s Now it becomes v = d/t 343m/s = d / 3.00)

OpenStudy (anonymous):

@roadjester

OpenStudy (anonymous):

@kewlgeek555 @johnweldon1993

OpenStudy (anonymous):